# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@
Newton numbers, vanishing polytopes and algebraic degrees
Fedor Selyanin
Consider a polynomial $f$ with a convenient Newton polytope $P$ and generic complex coefficients. By the global version of the Kouchnirenko formula, the hypersurface $\{f=0\}\subset\mathbb{C}^{n}$ has the homotopy type of a bouquet of $(n-1)$-spheres, and the number of spheres is given by a certain alternating sum of volumes, called the Newton number $\nu(P)$. Using the Furukawa-Ito classification of dual defective sets, we classify convenient Newton polytopes with vanishing Newton numbers as certain Cayley sums called $B_{k}$-polytopes. These $B_{k}$-polytopes generalize the $B_{1}$- and $B_{2}$-facets appearing in the local monodromy conjecture in the Newton non-degenerate case. Our classification provides a partial solution to the Arnold’s monotonicity problem.
@@
We generalize the Newton number in two independent ways: the $\ell$-Newton number and the $e$-Newton number. The $\ell$-Newton number comes from Ehrhart theory, namely, from certain generalizations of Katz-Stapledon decomposition formulas. It is the main ingredient in the proof of the thinness of the $B_{k}$-polytopes. The $e$-Newton number is the number of points of zero-dimensional critical complete intersections. Vanishing of the $e$-Newton number characterizes the dual defective sets. The $e$-Newton number calculates the algebraic degrees (Maximum Likelihood, Euclidean Distance, and Polar degrees). For instance, we show that all the known formulas for the algebraic degrees in the Newton non-degenerate case are implied by basic properties of the $e$-Newton number.
@@ 1 Introduction
Recall that the support of a polynomial $f\in\mathbb{C}[z_{1},\dots,z_{n}]$ is the set of the exponents of its non-zero monomials and the Newton polytope is the convex hull of the support. A convex polytope $P\subset\mathbb{R}^{n}_{\geq 0}$ is called convenient (cf. Definition 1.6) if it contains the origin and a segment on each coordinate axis.
@@
Consider a generic polynomial $f\in\mathbb{C}[z_{1},\dots,z_{n}]$ with Newton polytope $P$. By the global version of the Kouchnirenko theorem (see [Kou76] and [Bro88, Proposition 3.3]) the hypersurface $\{f=0\}\subset\mathbb{C}^{n}$ has the homotopy type of a bouquet of $\nu(A)$ of $(n-1)$-spheres, where the number $\nu(A)$ equals the number of critical points of $f$ in $\mathbb{C}^{n}$. Using the Furukawa-Ito classification of dual defective sets [FI21], we classify all negligible polytopes in $\mathbb{R}^{n}_{\geq 0}$. We give the answer in terms of certain Cayley sums (we recall them in §5.1) which we call $\mathbf{B_{k}}$-polytopes.
@@
where ‘‘$+$’’ is the Minkowski sum. Then the Cayley sum $P_{0}\ast P_{1}\ast\cdots\ast P_{k}\subset\mathbb{R}^{n-k}\oplus\mathbb{R}^{k}_{\geq 0}$ is called $\mathbf{B_{k}}$-polytope based on $\mathbb{R}^{k}_{\geq 0}$.
The notion of $B_{k}$-polytopes refers to the so-called $B_{1}$- and $B_{2}$-facets from [LV11] and [ELT22] arising in the context of the monodromy conjecture (see [Ve] for a general survey). That is, $B_{1}$- and $B_{2}$-polytopes are natural projections of the $B_{1}$- and $B_{2}$-facets (for instance, this projection is illustrated in [ELT22, Figure 2 on p.17]). The definitions of a convenient polytope and its Newton number can be extended to the case of the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ without significant changes (see Definitions 1.6 and 4.5).
@@
A convenient polytope in the cone $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ is negligible (i.e. its Newton number vanishes) if and only if it is a $B_{k}$-polytope based on a coordinate subspace $\mathbb{R}^{k}_{\geq 0}\subset\mathbb{R}^{m}_{\geq 0}$.
The latter theorem for all the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ is needed for the Arnold’s monotonicity problem (see [Arn, (1982-16)] for its initial formulation and e.g. [Sel24] for a more up to date version). In [Sel24, Theorem 3.3.(1)] the Arnold’s monotonicity problem in dimension $n+1$ is reduced to the classification of (not necessarily convex) convenient lattice polytopes with vanishing Newton number in the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. Thus, the $B_{k}$-theorem for all the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ represents a significant advance in the Arnold’s monotonicity problem.
In what follows, we define two types of Newton numbers: the $\mathbf{\ell}$-Newton number and the $\mathbf{e}$-Newton number. The $\ell$-Newton number arises from generalizations of Katz-Stapledon formulas (from Ehrhart theory). The $\ell$-Newton number is essential for proving that the $B_{k}$-polytopes are thin. We also believe that the $\ell$-Newton number is of independent combinatorial interest and that it admits a natural geometric interpretation.
The $e$-Newton number counts critical points of generic polynomials with given support. Its vanishing provides a criterion for dual defective sets. It counts algebraic degrees in the Newton non-degenerate case. All the known formulas for the algebraic degrees in the Newton non-degenerate case (maximum likelihood degree [Kh78], [CHKS06], [LNRW23], Euclidean distance degree [BSW22], [TT24], polar degree [Huh13]) follow from basic properties of the $e$-Newton number.
Unlike the usual Newton number (see [GKZ94, Example 2.5, p.353]) they are non-negative for arbitrary cones and convenient polytopes in it. Both $\ell$- and $e$-Newton numbers coincide with the usual Newton number in the case of convenient polytopes in $\mathbb{R}^{n}_{\geq 0}$. Consider a convenient polytope in $P\subset\mathbb{R}^{n}_{\geq 0}$ and a generic polynomial $f$ with Newton polytope $P$. Then the number of critical points of $f$ in $\mathbb{C}^{n}$ provides the calculation (similar to [Oka79, Section 4]) of the $e$-Newton number, and counting the middle homology groups of $\{f=0\}$ corresponds to the $\ell$-Newton number since the $\ell$-Newton number is consistent with the Stapledon’s formula for monodromy from [Stap17] (see Example 4.8).
@@
Does the $\ell$-Newton number of convenient polytopes in arbitrary cones have a natural geometric meaning?
The $e$-Newton number and the $\ell$-Newton number are generally different. In §5.3 we compare them with the usual Newton number. We conjecture that the $\ell$-Newton number is greater than or equal to the $e$-Newton number (see Conjecture 5.28).
1.1 $\ell$-Newton number and thin polytopes
Ehrhart theory studies lattice points in lattice polytopes. The $h^{*}$-polynomial of lattice polytopes is one of the main objects in Ehrhart theory. The local $h^{*}$-polynomial (or $\ell$-polynomial) is a natural refinement of the $h^{*}$-polynomial. Katz-Stapledon decomposition formulas provide decompositions of the $h^{*}$- and the local $h^{*}$-polynomials via polyhedral subdivision. These decomposition formulas also incorporate the $h$-polynomials of Eulerian posets and local $h$-polynomials (or $\ell$-polynomials) of strong formal subdivisions. We recall it all in §2.
Recall that ‘‘$F\leq C$’’ means ‘‘$F$ is a face of the cone $C$’’ (see Example 2.9).
@@ 1.1 $\ell$-Newton number and thin polytopes
A convex polytope $P\subset C$ in a cone $C$ is called convenient if for any $F\leq C$ we have $\dim(F\cap P)=\dim(F)$. The boundary $\partial_{C}(P)$ is the boundary of $P$ in the topology of $C$.
@@
Thus, by Theorem 1.3, the $\mathbf{B_{k}}$-polytopes are thin.
@@
This also motivates the following definition and a generalization of the Arnold’s monotonicity problem (solved for the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ in the $B_{k}$-Theorem 1.3).
@@
Let us say a few words concerning the agglutination technique from §3 and §4 which defines the $\ell_{\mathcal{C}}^{*}$-polynomial and the $\ell$-Newton number. It was initially developed to prove Corollary 1.7 but we believe that it is of independent interest. Remarkably, we do not know a proof of Corollary 1.7 avoiding this rather complicated technique. What is the most surprising is that our proof of Corollary 1.7 crucially relies on the unimodality of the local $h$-polynomial of regular polyhedral subdivisions (see [KS16, Theorem 6.1]).
@@
Is there a simpler proof of Corollary 1.7?
@@
We generalize the definitions of the $h^{*}$- and the local $h^{*}$-polynomials to the case of arbitrary strong formal subdivisions whose elements in the source are lattice polytopes (we call them polyhedral strong formal subdivisions). The Katz-Stapledon decomposition formulas are trivially extended to this case.
Our main examples of polyhedral strong formal subdivisions is the agglutination of a convenient polytope $P$ in a cone $C$ with itself along the corresponding boundary $\partial_{C}(P)$. We denote the corresponding $h^{*}$- and local $h^{*}$-polynomials by $h^{*}_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\mathcal{P};t)$ and $\ell^{*}_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\mathcal{P};t)$ respectively. The $h^{*}_{\mathcal{C}}$- and $\ell_{\mathcal{C}}^{*}$-polynomials of agglutinated polytopes have nonnegative coefficients just as the usual ones. The $\ell$-Newton number is the evaluation $\frac{1}{2}\ell^{*}_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\mathcal{P};1)$ which is a non-negative integer for every cone $C$ and any convenient polytope $P$ in it.
1.2 $e$-Newton number and Algebraic degrees
Consider a finite lattice set $\mathbf{P}\subset\mathbb{Z}^{n}$ and a point $X$. If $X$ is in $\mathbb{Z}^{n}$ then the $\mathbf{e}$-Newton number $\nu^{e}_{X}(\mathbf{P})$ is the number of critical points of a generic polynomial with support $\mathbf{P}-X$ (i.e. $\mathbf{P}$ shifted by the vector $-X$) in the torus $(\mathbb{C}\setminus 0)^{n}$. It can be generalized for an arbitrary point $X\in\mathbb{Q}\mathbb{P}^{n}$ as the number of points of a zero-dimensional critical complete intersection introduced in [Es17, Section 2.7]. We recall this together with the combinatorial formulas in §5.2.
All results concerning the $e$-Newton numbers discussed in the paper boil down to [Es17, Lemma 2.50] together with the expression for Euler obstructions of toric varieties [MT11] (see [Es24b, Example 1.7(6)] and [Es24b, Theorem 1.8] for more details). Many important special cases of this formula for the $e$-Newton number were rediscovered independently by other authors (e.g. see [FS24, Theorem A.; Theorem 3.8], where the theorems can be rewritten in terms of the $e$-Newton numbers and this rewriting immediately straightens and generalizes the results). We believe that our work will promote the zero-dimensional critical complete intersections.
Formulas for the algebraic degrees in the Newton non-degenerate case (maximum likelihood degree [CHKS06], [LNRW23], Euclidean distance degree [BSW22], [TT24], polar degree [Huh13]) boil down to some basic properties of the $e$-Newton numbers. We show this in detail in §5.4 and refine the results concerning the algebraic degrees. Comparing these with the $\ell$-Newton number, we, for instance, obtain the following unexpected corollary.

### v2

@@
Abstract
Consider a polynomial $f$ with a convenient Newton polytope $P$ and generic complex coefficients. By the global version of the Kouchnirenko formula, the hypersurface $\{f=0\}\subset\mathbb{C}^{n}$ has the homotopy type of a bouquet of (n1)(n-1)-spheres, and the number of spheres is given by a certain alternating sum of volumes, called the Newton number $\nu(P)$. Using the Furukawa-Ito classification of dual defective sets, we classify convenient Newton polytopes with vanishing Newton numbers as certain Cayley sums called $B_{k}$-polytopes. These $B_{k}$-polytopes generalize the $B_{1}$- and $B_{2}$-facets appearing in the local monodromy conjecture in the Newton non-degenerate case. Our classification provides a partial solution to Arnold’s monotonicity problem.
@@ Abstract
We generalize the Newton number in two independent ways: the $\ell$-Newton number and the $e$-Newton number. The $\ell$-Newton number comes from Ehrhart theory, namely, from certain generalizations of Katz-Stapledon decomposition formulas, and its properties are central to our proof that the $B_{k}$-polytopes are thin. The $e$-Newton number is the number of points of zero-dimensional critical complete intersections. Vanishing of the $e$-Newton number characterizes dual defective sets. Furthermore, the $e$-Newton number calculates algebraic degrees (such as Maximum Likelihood, Euclidean Distance and Polar degrees). For instance, we show that all known formulas for these algebraic degrees in the Newton non-degenerate case are implied by basic properties of the $e$-Newton number.
@@ 1 Introduction
Recall that the support of a polynomial $f\in\mathbb{C}[z_{1},\dots,z_{n}]$ is the set of the exponents of its nonzero monomials and the Newton polytope is the convex hull of the support. A convex polytope $P\subset\mathbb{R}^{n}_{\geq 0}$ is called convenient (cf. Definition 1.6) if it contains the origin and a segment on each coordinate axis.
@@
Consider a generic polynomial $f\in\mathbb{C}[z_{1},\dots,z_{n}]$ with Newton polytope $P$. By the global version of the Kouchnirenko theorem (see [Kou76] and [Bro88, Proposition 3.3]) the hypersurface $\{f=0\}\subset\mathbb{C}^{n}$ has the homotopy type of a bouquet of $\nu(P)$ of (n1)(n-1)-spheres, and $\nu(P)$ also equals the number of critical points of $f$ in $\mathbb{C}^{n}$. Using the Furukawa-Ito classification of dual defective sets ([FI21]; see Theorem 6.5), we classify all negligible polytopes in $\mathbb{R}^{n}_{\geq 0}$. We give the answer in terms of Cayley sums (we recall them in §5.1) which we call $\mathbf{B_{k}}$-polytopes.
@@
where ‘‘++’’ is the Minkowski sum. Then the Cayley sum $P_{0}\ast P_{1}\ast\cdots\ast P_{k}\subset\mathbb{R}^{n-k}\oplus\mathbb{R}^{k}_{\geq 0}$ is called $\mathbf{B_{k}}$-polytope based on $\mathbb{R}^{k}_{\geq 0}$.
The notion of $B_{k}$-polytopes refers to the so-called $B_{1}$- and $B_{2}$-facets from [LV11] and [ELT22] arising in the context of the monodromy conjecture (see [Ve] for a general survey). That is, $B_{1}$- and $B_{2}$-polytopes are natural projections of the $B_{1}$- and $B_{2}$-facets (for instance, this projection is illustrated in [ELT22, Figure 2 on p.17]). Note that the $B_{k}$-polytopes differ from the $B$-polytopes [ELT22, Definition 1.4], for more details see §6.3.1.
The definitions of a convenient polytope and its Newton number can be extended to the case of the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ without significant changes (see Definitions 1.6 and 4.5).
@@
A convenient polytope $P$ in the cone $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ is negligible (i.e. $\nu(P)=0$) if and only if it is a $B_{k}$-polytope based on a coordinate subspace $\mathbb{R}^{k}_{\geq 0}\subset\mathbb{R}^{m}_{\geq 0}$.
The latter theorem for all the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ is needed for Arnold’s monotonicity problem. Arnold’s monotonicity problem suggests to study the (strict) monotonicity conditions of the Newton number (see [Arn, (1982-16)] for its initial formulation and e.g. [Sel24] for a more up to date version). In [Sel24, Theorem 3.3.(1)] Arnold’s monotonicity problem in dimension $n+1$ is reduced to the classification of (not necessarily convex) convenient lattice polytopes with vanishing Newton number in the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. Thus, the $B_{k}$-theorem for all the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ represents a significant advance in Arnold’s monotonicity problem. A wrong answer to Arnold’s monotonicity problem was conjectured in the end of the introduction of [BKW19] and claimed in [L-AMS20, Theorem 2.25 and Remark 2.22]. For the simplest counterexample (a Morse function in dimension $4$), see [Sel24, Example 1.1]. Also note that using [Sel24, Theorem 3.3.(1)] one can obtain a counterexample from every $B_{k}$-polytope, which is not a $B_{1}$-polytope.
In what follows, we define two types of Newton numbers: the $\mathbf{\ell}$-Newton number and the $\mathbf{e}$-Newton number. The $\ell$-Newton number arises from generalizations of Katz-Stapledon formulas (from Ehrhart theory). The $\ell$-Newton number is the main tool for proving that the $B_{k}$-polytopes are thin. We also believe that the $\ell$-Newton number is of independent combinatorial interest and that it admits a natural geometric interpretation.
The $e$-Newton number counts critical points in the complex torus of generic polynomials with given support. Its vanishing provides a criterion for dual defective sets. It counts algebraic degrees in the Newton non-degenerate case. All the known formulas for the algebraic degrees in the Newton non-degenerate case (maximum likelihood degree [Kh78], [CHKS06], [LNRW23], Euclidean distance degree [BSW22], [TT24], polar degree [Huh13]) follow from basic properties of the $e$-Newton number.
Unlike the usual Newton number (see [GKZ94, Example 2.5 on p.353]) our Newton numbers are non-negative for arbitrary cones and convenient polytopes in it. Both $\ell$- and $e$-Newton numbers coincide with the usual Newton number in the case of convenient polytopes in $\mathbb{R}^{n}_{\geq 0}$. Consider a convenient polytope in $P\subset\mathbb{R}^{n}_{\geq 0}$ and a generic polynomial $f$ with Newton polytope $P$. Then the number of critical points of $f$ in $\mathbb{C}^{n}$ provides the calculation (similar to [Oka79, Section 4] in the local case) of the $e$-Newton number, and counting the middle homology groups of $\{f=0\}$ corresponds to the $\ell$-Newton number since the $\ell$-Newton number is consistent with the Stapledon’s formula for monodromy from [Stap17] (see Example 4.8).
@@
Does the $\ell$-Newton number of convenient polytopes in arbitrary cones have a natural geometric interpretation?
The $e$-Newton number and the $\ell$-Newton number are generally different. In §5.3 we compare them with the usual Newton number. We believe that the $\ell$-Newton number is greater than or equal to the $e$-Newton number (see Conjecture 5.29).
1.1 $\mathbf{\ell}$-Newton number and thin polytopes
Ehrhart theory generalizes Pick’s theorem by studying the number of lattice points in integer dilations of lattice polytopes. The $h^{*}$-polynomial of lattice polytopes is a convenient way to encode this data. The local $h^{*}$-polynomial (or $\ell^{*}$-polynomial) is a natural refinement of the $h^{*}$-polynomial. Katz-Stapledon decomposition formulas provide decompositions of the $h^{*}$- and the local $h^{*}$-polynomials for polyhedral subdivision. These decomposition formulas also incorporate the $h$-polynomials of Eulerian posets and local $h$-polynomials (or $\ell$-polynomials) of strong formal subdivisions.We recall all of these concepts in §2.
Recall that $F\leq C$’’ means ‘‘$F$ is a face of the cone $C$’’ (see Example 2.9).
@@ 1.1 $\mathbf{\ell}$-Newton number and thin polytopes
A convex polytope $P\subset C$ is called convenient (in the cone $C$) if for every face $F\leq C$ we have $\dim(F\cap P)=\dim F$. The boundary $\partial_{C}(P)$ is the boundary of $P$ in the topology of $C$.
@@
Thus, for a $B_{k}$-polytope $P$, Theorem 1.3 gives $\nu(P)=0$, which implies $\ell^{*}(P;1)=0$. Therefore, $\mathbf{B_{k}}$-polytopes are thin.
@@
This also motivates the following definition and a generalization of Arnold’s monotonicity problem (solved for the cones $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ in the $B_{k}$-Theorem 1.3).
@@
We now briefly describe the agglutination technique from §3 and §4, which is used to define the $\ell_{\mathcal{C}}^{*}$-polynomial and the $\ell$-Newton number. It was initially developed to prove Corollary 1.7 but we believe that it is of independent interest. Remarkably, we do not know a proof of Corollary 1.7 avoiding this rather complicated technique. Most surprisingly, our proof of Corollary 1.7 crucially relies on the unimodality of the local $h$-polynomial of regular polyhedral subdivisions (see [KS16, Theorem 6.1]).
@@
Is there a simpler proof of Corollary 1.7? Is there a simpler way to verify that $B_{k}$-polytopes are thin?
@@
We generalize the definitions of the $h^{*}$- and the local $h^{*}$-polynomials to the case of arbitrary strong formal subdivisions whose elements in the source are lattice polytopes (we call them polyhedral strong formal subdivisions). Katz-Stapledon decomposition formulas are trivially extended to this case.
Our main example of polyhedral strong formal subdivisions is the agglutination of a convenient polytope $P$ in a cone $C$ with itself along the corresponding boundary $\partial_{C}(P)$ of $P$ in the topology of the cone $C$. We denote the corresponding $h^{*}$- and local $h^{*}$-polynomials by $h^{*}_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P};t)$ and $\ell^{*}_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P};t)$, respectively. The $h^{*}_{\mathcal{C}}$- and $\ell_{\mathcal{C}}^{*}$-polynomials of agglutinated polytopes have nonnegative coefficients, mirroring the property of the standard $h^{*}$- and $\ell^{*}$-polynomials. The $\ell$-Newton number is the evaluation $\frac{1}{2}\ell^{*}_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P};1)$ which is a non-negative integer for every cone $C$ and any convenient polytope $P$ in it.
1.2 $\mathbf{e}$-Newton number and Algebraic degrees
Consider a finite lattice set $\mathbf{P}\subset\mathbb{Z}^{n}$ and a point $X$. If $X$ is in $\mathbb{Z}^{n}$ then the $\mathbf{e}$-Newton number $\nu^{e}_{X}(\mathbf{P})$ is the number of critical points of a generic polynomial with support $\mathbf{P}-X$ (i.e. $\mathbf{P}$ shifted by the vector X-X) in the torus $(\mathbb{C}\setminus 0)^{n}$. It can be generalized for an arbitrary point $X\in\mathbb{Q}\mathbb{P}^{n}$ as the number of points of a zero-dimensional critical complete intersection introduced in [Es18, Section 2.7]. We recall this together with the combinatorial formulas in §5.2.
All results concerning the $e$-Newton numbers discussed in the paper boil down to [Es18, Lemma 2.50] together with the expression for Euler obstructions of toric varieties [MT11] (see [Es24b, Example 1.7(6)] and [Es24b, Theorem 1.8] for more details). Many important special cases of this formula for the $e$-Newton number were rediscovered independently by other authors (e.g., see [FS24, Theorem A.; Theorem 3.8], where the theorems can be rewritten in terms of the $e$-Newton numbers and this rewriting immediately simplifies and generalizes the results, see Remark 5.39). We believe that our work will help to popularize the theory of (zero-dimensional) critical complete intersections.
Formulas for the algebraic degrees in the Newton non-degenerate case (maximum likelihood degree [CHKS06], [LNRW23], Euclidean distance degree [BSW22], [TT24], polar degree [Huh13]) boil down to some basic properties of the $e$-Newton numbers. We show this in detail in §5.5 and refine the results concerning the algebraic degrees. The $e$-Newton number also computes the support function of the Newton polytope of the $\mathbf{A}$-discriminant from [GKZ94], see §5.4. Comparing these with the $\ell$-Newton number, we, for instance, obtain the following unexpected corollary.